OpenStudy (anonymous):

At the candy store a soda and a pretzel cost $1.70. What is the cost of each if the soda cost $1.00 more than the pretzel?

OpenStudy (anonymous):

Take inequality in use

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\(x\) is the price of the soda. \[x + (x - 1.00) = 1.70\]

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

\(x - 1.00\) is the price of the pretzel, since the pretzel is $1 less than the price of the soda (which, you'll recall, is \(x\)).

OpenStudy (anonymous):

This is an algebraic problem. Can you solve the equation I typed above?

OpenStudy (aripotta):

p + s = 1.70 s = p + 1.00 using the substitution method: p + p + 1.00 = 1.70 can you solve for p now?

OpenStudy (anonymous):

wats the answer

OpenStudy (aripotta):

i actually gave you the answer in the chat, but i don't think you were paying attention. so now you have to work it out >:)

OpenStudy (anonymous):

no u didnt

OpenStudy (aripotta):

well, i just deleted it :l this isn't hard, i'm trying to help

OpenStudy (aripotta):

what's the first step in solving for p? p + p + 1.00 = 1.70

OpenStudy (anonymous):

wats the answer